4. Show that n2 + 3n is divisible by 2 for all natural numbers n 21
Answer #1

penian²+ 3n Trie.-tar nal p11)=(0)²+3610 =1+3 ..P(1) 24 ii PCD is true. for n=1. Assume true for nak. p(l) =) (C²+316.agm MEN2 for is divisible by all natural numbers nzi, Pelcro)